Kohler Corporation reports the following components of stockholders’ equity at December 31, 2018.

Common stock—$15 par value, 100,000 shares authorized,
55,000 shares issued and outstanding
$ 825,000
Paid-in capital in excess of par value, common stock 70,000
Retained earnings 400,000
Total stockholders' equity $ 1,295,000


During 2019, the following transactions affected its stockholders’ equity accounts.

Jan. 2 Purchased 4,500 shares of its own stock at $25 cash per share.
Jan. 5 Directors declared a $4 per share cash dividend payable on February 28 to the February 5 stockholders of record.
Feb. 28 Paid the dividend declared on January 5.
July 6 Sold 1,688 of its treasury shares at $29 cash per share.
Aug. 22 Sold 2,812 of its treasury shares at $22 cash per share.
Sept. 5 Directors declared a $4 per share cash dividend payable on October 28 to the September 25 stockholders of record.
Oct. 28 Paid the dividend declared on September 5.
Dec. 31 Closed the $388,000 credit balance (from net income) in the Income Summary account to Retained Earnings.


Required:

1. Prepare journal entries to record each of these transactions.
2. Prepare a statement of retained earnings for the year ended December 31, 2019.
3. Prepare the stockholders’ equity section of the company’s balance sheet as of December 31, 2019.

Prepare a statement of retained earnings for the year ended December 31, 2019. (Amounts to be deducted should be indicated by a minus sign.)

KOHLER CORPORATION
Statement of Retained Earnings
For Year Ended December 31, 2019
0
$0
  • Required 1
  • Prepare the stockholders’ equity section of the company’s balance sheet as of December 31, 2019.

    KOHLER CORPORATION
    Stockholders' Equity Section of the Balance Sheet
    December 31, 2019
    Total contributed capital 0
    Total stockholders' equity $0
Answer #1
1
Date General Journal Debit Credit
2-Jan Treasury stock, Common 112,500 =4500*25
Cash 112,500
5-Jan Retained earnings 202,000 =(55000-4500)*4
Common dividend payable 202,000
28-Feb Common dividend payable 202,000
Cash 202,000
6-Jul Cash 48,952 =1688*29
Treasury stock, Common 42,200 =1688*25
Paid-in capital, Treasury stock 6,752
22-Aug Cash 61,864 =2812*22
Paid-in capital, Treasury stock 6,752
Retained earnings 1,684
Treasury stock, Common 70,300 =2812*25
5-Sep Retained earnings 220,000 =55000*4
Common dividend payable 220,000
28-Oct Common dividend payable 220,000
Cash 220,000
31-Dec Income summary 388,000
Retained earnings 388,000
2
KOHLER CORPORATION
Statement of Retained Earnings
For Year Ended December 31, 2019
Retained earnings, Dec. 31, 2018 400,000
Add: Net income 388,000
788,000
Less: Cash dividends declared -422,000
Less: Treasury stock reissuances -1,684
Retained earnings, Dec. 31, 2019 364,316
3
KOHLER CORPORATION
Stockholders' Equity Section of the Balance Sheet
December 31, 2019
Common stock - $15 par value 825,000
Paid-in capital in excess of par value, common stock 70,000
Total contributed capital 895,000
Retained earnings 364,316
Total stockholders' equity 1,259,316